OpenStudy (anonymous):

Please help ? Using complete sentences, explain the two ways to simplify sqrt 147 and provide the simplified form.

OpenStudy (raden):

the sum all digits 1+4+7=12 this number divisible by 3, because 12 divisible by 3 now, 147/3 = 49 so, sqrt(147) can rewriten as sqrt(3 * 49) = sqrt(3) * sqrt(49) = sqrt(3) * 7 = 7 sqrt(3)

OpenStudy (anonymous):

thank you so much :)

OpenStudy (raden):

you're welcome :)